# OLATverse: A Large-scale Real-world Object Dataset with Precise Lighting Control

### v1

@@ 1 Introduction
The appearance of real-world objects is the result of the complex interaction between geometry, material, and lighting conditions. Acquiring high-fidelity material and geometry of objects, and synthesizing photorealistic appearance under novel illuminations, remain fundamental challenges in computer vision and computer graphics, with widespread applications in film and gaming industries, autonomous driving, robotics, and VR/AR. With the advances in deep learning and generative models, recent years have witnessed rapid progress in 3D generation and reconstruction [45, 51, 53, 55, 35, 40, 47], relighting [23, 60, 64, 59, 34] and inverse rendering [3, 66, 31, 16, 61, 30]. However, due to the lack of high-fidelity large-scale real-world datasets, the majority of these techniques are trained on synthetic datasets [9, 57] or evaluated on small-scale real datasets [3, 33]. As a result, the synthesized results are typically limited in the realism, and their performance in real-world scenarios cannot be reliably assessed because of the significant domain gap between real and synthetic data.
@@
Considerable effort has recently been devoted to constructing large-scale and high-fidelity object-centric datasets. However, due to complexity in hardware and data processing pipelines, the existing datasets remain constrained in at least one of three key aspects: quality [44, 7, 9], scale [21, 71, 27, 18, 50, 59, 28, 3, 11, 33] or precise lighting control [57, 12]. As summarized in Tab. 1, the first category of datasets [44, 7, 9] comprises synthetic and hybrid objects. While the scale of these datasets is large, they typically lack realism and exhibit significant variance in object quality. The second category of datasets [57, 12] contains thousand-scale real-world objects, offering both realism and scale. However, the material creation in these datasets heavily relies on manual annotation, and the limited illumination setups prevent accurate simulation of complex light transport. The third group of datasets [3, 33, 11] captures objects under precisely controlled illumination, but this group is typically small in both scale and diversity, restricting their applications in comprehensive benchmarking and generative models training. To the best of our knowledge, no existing object dataset simultaneously provides both large-scale coverage and high-fidelity appearance.
To address these challenges, we propose OLATverse, the first large-scale real dataset that provides high-fidelity images of a diverse set of objects captured under precisely-controlled illumination and camera configurations. Our dataset consists of $765$ real-world objects, covering a wide range of material categories (e.g., wood, stone, leather, plastic, metal, paper, plaster, fabric, ceramic) and $18.5\%$ LVIS categories [20] from common to uncommon objects. OLATverse is acquired using lightstage setup  [8], which enables One-Light-at-a-Time (OLAT) capture of real-world objects and provides rich information for analyzing the complex interaction between surface reflectance properties and light sources. Specifically, each object is captured using 35 well-calibrated DSLR cameras and illuminated by 331 individually controlled light sources, simulating a set of illuminations including uniform lightings, OLATs, gradient illumination and pre-defined environmental illuminations. In total, this setup yields over $9M$ high-fidelity images. In addition, we develop efficient semi-automatic mask processing pipeline to extract high-quality masks for each object captured under multiple views. Furthermore, we employ polarized gradient illumination to recover surface normal and diffuse albedo, following techniques proposed by Ma et al. [36]. These auxiliary data are particularly valuable for evaluating and supervising multi-modal tasks.
@@
ABO [7] 8K ENV 3
ShapeNet-Intr [44] 31K ENV 36
TexVerse [67] 818K
Objaverse [9] 858K
@@
Standford-ORB [28] 14 ENV 7 70 scanner+camera
DTC [12] 2K ENV 2 120 scanner+camera
OmniObject3D [57] 6K scanner
DiLiGenT-MV [71] 5 OLAT 96 20 scanner
GS3 [3] 6 OLAT Unspec. Unspec. lightstage
NRHint [59] 7 OLAT Unspec. Unspec. smartphones
DIR [6] 16 OLAT 144 2 LCD display
MIT-Intrinsic [18] 20 OLAT 10 1 camera
ReNe [50] 20 OLAT 40 50 robots
SSS-GS [11] 20 OLAT 167 158 lightstage
OpenIllumination [33] 64 PAT+OLAT 13+142 72 lightstage
@@ 2.1 Object-centric Datasets
The acquisition of high-fidelity and large-scale datasets is critical for advancing data-driven methods in inverse rendering and relighting tasks. Recent research has introduced different datasets that capture the geometry and reflectance properties of full-body human avatars [68, 48, 54, 49], human faces [43, 62, 41], planar material surfaces [10, 52, 37, 70, 69] and 3D objects [44, 7, 9, 21, 71, 27, 18, 50, 59, 57, 28, 12, 3, 11, 33, 67, 6]. In comparison, our work focuses on appearance capture of real-world objects. We categorize the existing object-centric datasets into several subgroups, as summarized in Tab. 1. The first group focuses on synthetic objects, which are highly scalable due to the procedural generation pipelines and the availability of public 3D models. ShapeNet-Intrinsics [44], for example, constructs a large-scale object intrinsic dataset from 3D models in the ShapeNet [4] collection. ABO [7] introduces synthetic household objects with complex geometry and materials. Objaverse [9] and TexVerse [67] significantly expand the diversity of objects by incorporating both synthetic and scanned real assets. While these datasets are large-scale, the quality of these datasets varies significantly across different objects, and a domain gap still remains between synthetic and real-world data. Consequently, effective curation strategies are required to filter out low-quality data, limiting their applicability in realistic 3D vision and relighting.
@@
Figure 2: (a) We visualize the statistics of OLATverse, including the material distribution and high-level object category distribution. (b) We also show comparison against OpenIllumination [33] for the six largest material and object categories in terms of object count.
@@
The second group of datasets focuses on collecting real-world objects with high-fidelity appearance. Such datasets are typically collected either in specialized studios or in-the-wild environments. Among this group, DTC [12] and OmniObject3D [57] provide the meshes of thousands of real objects via 3D scanning. Although these two datasets are large-scale, they offer limited support for appearance analysis and lack control over lighting conditions. Specifically, OmniObject3D does not support appearance capture, and DTC only contains only 50 objects captured under two environmental illuminations. In addition, the material maps provided in DTC are manually created by artists to approximate the realistic appearance, which introduces a gap to the real-world appearance.
@@
The third group of datasets [3, 33, 11] captures OLATs in studios with a well-designed lighting setups (e.g., lightstage [8], LCD, smartphone flash, etc.), facilitating detailed analysis of light transport. However, due to the complexity of the setup and capture procedure, these datasets remain small in scale and limited in diversity. The largest OLAT dataset, OpenIllumination [33], contains only $64$ objects, limiting its application for comprehensive benchmarking and generative model training. In comparison with existing object-centric datasets, OLATverse encompasses $765$ real-world objects spanning diverse material categories and provides high-fidelity appearance under precisely controlled lighting, thereby achieving a combination of scale, diversity, and realism.
@@ 2.2 Inverse Rendering &\& Relighting
The goal of inverse rendering is to recover the intrinsic properties of scenes or objects (e.g., albedo, normal, roughness, and metallicity) to reproduce plausible renderings under novel illuminations. Traditional approaches typically leverage advanced differentiable renderers [39, 42, 29] under reliable constraints provided by geometry to jointly optimize intrinsic properties and lighting by minimizing the discrepancy between input and rendered images. With the advances in deep learning, Neural Radiance Fields (NeRF)[38] is introduced to jointly encode geometry and appearance in a single multi-layer perceptron (MLP), producing photorealistic rendering via volume rendering. NeRF demonstrates superior performance but suffers from slow rendering speed. To address this limitation, 3D Gaussian Splatting (3DGS)[24] is proposed to represent scenes with localized Gaussian kernels and can offer real-time efficiency through an efficient differentiable splatting process. Leveraging the benefits of these two representations, many recent inverse rendering methods [66, 46, 16, 31, 13, 63, 34, 22, 14] extend NeRF or 3DGS with additional optimizable Bidirectional Reflectance Distribution Function (BRDF) modules and a differentiable renderer to recover intrinsic properties and lighting from multi-view input images. More recently, with the rapid progress of diffusion-based generative models, a new branch of research leverages the strong priors of pretrained diffusion models for inverse rendering [61, 58, 30, 5] and relighting [23, 60] tasks. These methods typically fine-tune pretrained diffusion models on datasets annotated with material maps or renderings under novel illuminations. During inference, the diffusion model extracts material reflectance properties or relit images through inverse denoising steps. In this work, we demonstrate that OLATverse serves as a comprehensive benchmark for object-centric inverse rendering and relighting tasks, and we evaluate several representative approaches as baselines on our selected validation dataset.
@@ 3.1 Overview
OLATverse comprises $765$ real-world objects with high diversity in material types, object categories, and physical sizes. As shown in (a) of Fig. 2, it covers over $13$ material categories, including wood, stone, plaster, fabric, plastic, metal, food, plant, ceramics, leather, wax, rubber, and paper. Furthermore, OLATverse consists of a wide range of common and uncommon real-world objects, spanning over $18.5\%$ of LVIS categories [20], substantially outperforming existing real-world object datasets, such as OmniObject3D ($10.8\%$[57], OpenIllumination ($4\sim 5\%$[33] and DTC ($3\%$[12]. Additionally, OLATverse is not limited to object physical dimension, encompassing a wide range of sizes from $5\mathrm{cm}$ to $100\mathrm{cm}$, which is broader than OpenIllumination dataset ($10\sim 20\mathrm{cm}$). These attributes highlight the diversity and scale of OLATverse, establishing it as the first comprehensive OLAT dataset of real-world objects.
@@
We compare OLATverse against one representative OLAT real dataset, OpenIllumination [33]. We visualize the number of objects in the six largest material and high-level object categories, as illustrated in Fig. 2 (b). While OpenIllumination covers the similar material and object categories as OLATverse, the scale of each category is significantly smaller than our dataset. In addition, OpenIllumination does not provide auxiliary resources such as surface normals and diffuse albedo, limiting its applications in the multi-modal tasks.
@@ 3.3 Data Processing
Acquiring accurate camera parameters consistently across large-scale $765$ objects in a studio setup is a non-trivial task. Directly applying feature-based calibration algorithms to different objects leads to unstable and ambiguous calibration quality since the real-world objects vary widely in physical dimension, texture richness, and material properties, which limit the accuracy and consistency of the camera calibration process. To address this challenge, we leverage the fixed camera configuration of the lightstage, and perform a calibration session every $20\sim 30$ regular capture sessions, demonstrated in (b) of Fig. 3. During calibration session, we lock down several fixed objects with rich texture and Lambertian surfaces as calibration references, and employ the feature-based algorithms implemented in Metashape [1] to recover both intrinsic and extrinsic camera parameters. For subsequent capture sessions, the previously estimated camera parameters are reused. All calibration captures are conducted under uniform white illumination to ensure robust and consistent feature detection. We quantitatively evaluate the calibration accuracy by computing the mean reprojection error of triangulated keypoints across the reference objects, yielding an average error of $0.86$ pixels. The positions of individual lights are also provided within the same canonical system by measuring their physical location. We visualize the calibrated cameras, light sources, and the reconstructed mesh of one calibration reference object in  Fig. 5. As is shown, the reconstructed mesh qualitatively aligns with real captured frames, demonstrating the accuracy of calibration process.
@@
Unlike human-centric datasets [49], where Sapiens [25] is tailored specifically for human body segmentation, efficiently producing high-quality masks for large-scale objects in our capture setup remains a challenging task. A straightforward solution is to employ Segment-Anything (SAM) [26] with multiple bounding-box prompts for instance segmentation, similar to OpenIllumination [33]. However, this strategy is inefficient and difficult to scale for a large-scale dataset. To efficiently tackle this challenge, we develop a simple yet effective semi-automatic segmentation pipeline that combines background matting (bgMatting) [32], SAM [26] and RMBG-2.0 [2]. Specifically, we capture a foreground image $\mathbf{I}_{\text{fg}}$, containing the object and stand, and a background image $\mathbf{I}_{\text{bg}}$ only for the stand, as is shown in (d) of Fig. 3. We empirically observe that bgMatting is capable of roughly separating the object from the stand, but fails to preserve fine details around the contour of objects. SAM, when guided by bounding boxes generated from bgMatting, can achieve a clean separation between the stand and the object but at the cost of mask quality. In comparison to the other two methods, RMBG-2.0 captures the most detailed contours but consistently treats the stand as part of the foreground, yielding a wrong object mask. Therefore, we leverage the strengths of all three segmentation strategies, and compute the final object mask $\mathbf{M}_{\text{obj}}$ as follows:
@@
Figure 7: We visualize the inverse rendering and novel view synthesis results of several baseline methods (Mitsuba [39], GS3 [3],BiGS [34] and RNG [14]) evaluated on our validation dataset. In this figure, we show relit objects from inference views and light directions.
@@
where $\mathbf{M}_{1}$, $\mathbf{M}_{2}$ and $\mathbf{M}_{3}$ represent bgMatting [32], SAM [26] and RMBG-2.0 [2], respectively. $\mathbf{M}_{\text{stup}}$ represents the mask of the stand. In Eq. 1, case a is applied to lower camera views, where part of the objects may be occluded by the stand, while case b is used for all other views. Then we refine the intermediate results $\mathbf{M}_{\text{obj}}^{*}$ using morphological transformations and remove the disconnected components to obtain the final clean mask $\mathbf{M}_{\text{obj}}$. This pipeline achieves a success rate of $95\%$ across all objects from all views in our setup. For the remaining failure cases, we create a lightweight user interface for manual correction. In  Fig. 6, we illustrate masks generated by three individual segmentation strategies and our method.
@@
In addition to diverse illuminations, OLATverse provides pseudo ground truth surface normals and diffuse albedo as auxiliary data to support multimodal tasks. We recover surface normals through photometric stereo solutions [19] by analyzing radiance variations of images captured under gradient illuminations. For all cameras, we captured images $\mathbf{I}_{\text{cg}}$ under color gradient illumination. To further cancel out the view-dependent specular reflection on the non-Lambertian surfaces [36, 17], we apply linear polarized filters to five fixed cameras, and capture images under polarized gradient full bright illuminations, denoted as $\mathbf{I}_{\perp}$. The diffuse albedo $\mathbf{D}$ and surface normals $\mathbf{N}$ are computed as follows:

### v2

@@ 1 Introduction
The appearance of real-world objects is the result of the complex interaction between geometry, material, and lighting conditions. Acquiring high-fidelity material and geometry of objects, and synthesizing photorealistic appearance under novel illuminations, remain fundamental challenges in computer vision and computer graphics, with widespread applications in film and gaming industries, autonomous driving, robotics, and VR/AR. With the advances in deep learning and generative models, recent years have witnessed rapid progress in 3D generation and reconstruction [shimvdream, toussaint2025probesdf, wang2021neus, wang2023neus2, long2024wonder3d, poole2022dreamfusion, siddiqui2024meta], relighting [jin2024neural, zeng2024dilightnet, zhang2025scaling, zeng2023relighting, liu2024bigs] and inverse rendering [bi2024gs3, zhang2021nerfactor, liang2024gs, gao2024relightable, zeng2024rgb, liang2025diffusion]. However, due to the lack of high-fidelity large-scale real-world datasets, the majority of these techniques are trained on synthetic datasets [deitke2023objaverse, wu2023omniobject3d] or evaluated on small-scale real datasets [bi2024gs3, liu2023openillumination]. As a result, the synthesized results are typically limited in the realism, and their performance in real-world scenarios cannot be reliably assessed because of the significant domain gap between real and synthetic data.
@@
Considerable effort has recently been devoted to constructing large-scale and high-fidelity object-centric datasets. However, due to complexity in hardware and data processing pipelines, the existing datasets remain constrained in at least one of three key aspects: quality [shi2017learning, collins2022abo, deitke2023objaverse], scale [jensen2014large, zhou2013multi, kuang2022neroic, grosse2009ground, toschi2023relight, zeng2023relighting, kuang2023stanford, bi2024gs3, dihlmann2024subsurface, liu2023openillumination] or precise lighting control [wu2023omniobject3d, dong2025digital]. As summarized in Tab. 1, the first category of datasets [shi2017learning, collins2022abo, deitke2023objaverse] comprises synthetic and hybrid objects. While the scale of these datasets is large, they typically lack realism and exhibit significant variance in object quality. The second category of datasets [wu2023omniobject3d, dong2025digital] contains thousand-scale real-world objects, offering both realism and scale. However, the material creation in these datasets heavily relies on manual annotation, and the limited illumination setups prevent accurate simulation of complex light transport. The third group of datasets [bi2024gs3, liu2023openillumination, dihlmann2024subsurface] captures objects under precisely controlled illumination, but this group is typically small in both scale and diversity, restricting their applications in comprehensive benchmarking and generative models training. To the best of our knowledge, no existing object dataset simultaneously provides both large-scale coverage and high-fidelity appearance.
To address these challenges, we propose OLATverse, the first large-scale real dataset that provides high-fidelity images of a diverse set of objects captured under precisely-controlled illumination and camera configurations. Our dataset consists of $765$ real-world objects, covering a wide range of material categories (e.g., wood, stone, leather, plastic, metal, paper, plaster, fabric, ceramic) and $18.5\%$ LVIS categories [gupta2019lvis] from common to uncommon objects. OLATverse is acquired using lightstage setup  [debevec2000acquiring], which enables One-Light-at-a-Time (OLAT) capture of real-world objects and provides rich information for analyzing the complex interaction between surface reflectance properties and light sources. Specifically, each object is captured using 35 well-calibrated DSLR cameras and illuminated by 331 individually controlled light sources, simulating a set of illuminations including uniform lightings, OLATs, gradient illumination and pre-defined environmental illuminations. In total, this setup yields over $9M$ high-fidelity images. In addition, we develop efficient semi-automatic mask processing pipeline to extract high-quality masks for each object captured under multiple views. Furthermore, we employ polarized gradient illumination to recover surface normal and diffuse albedo, following techniques proposed by Ma et al. [ma2007rapid]. These auxiliary data are particularly valuable for evaluating and supervising multi-modal tasks.
@@
ABO [collins2022abo] 8K ENV 3
ShapeNet-Intr [shi2017learning] 31K ENV 36
Objaverse [deitke2023objaverse] 818K
TexVerse [zhang2025texverse] 858K
@@
Standford-ORB [kuang2023stanford] 14 ENV 7 70 scanner+camera
DTC [dong2025digital] 2K ENV 2 120 scanner+camera
OmniObject3D [wu2023omniobject3d] 6K scanner
DiLiGenT-MV [zhou2013multi] 5 OLAT 96 20 scanner
GS3 [bi2024gs3] 6 OLAT Unspec. Unspec. lightstage
NRHint [zeng2023relighting] 7 OLAT Unspec. Unspec. smartphones
DIR [choi2025real] 16 OLAT 144 2 LCD display
MIT-Intrinsic [grosse2009ground] 20 OLAT 10 1 camera
ReNe [toschi2023relight] 20 OLAT 40 50 robots
SSS-GS [dihlmann2024subsurface] 20 OLAT 167 158 lightstage
OpenIllumination [liu2023openillumination] 64 PAT+OLAT 13+142 72 lightstage
@@ 2.1 Object-centric Datasets
The acquisition of high-fidelity and large-scale datasets is critical for advancing data-driven methods in inverse rendering and relighting tasks. Recent research has introduced different datasets that capture the geometry and reflectance properties of full-body human avatars [zhou2023relightable, stratou2011effect, wang2025relightable, teufel2025humanolat], human faces [saito2024relightable, zhang2021neural, prao20253dpr], planar material surfaces [deschaintre2018single, vecchio2024matsynth, ma2023opensvbrdf, zhou2023photomat, zhou2021adversarial] and 3D objects [shi2017learning, collins2022abo, deitke2023objaverse, jensen2014large, zhou2013multi, kuang2022neroic, grosse2009ground, toschi2023relight, zeng2023relighting, wu2023omniobject3d, kuang2023stanford, dong2025digital, bi2024gs3, dihlmann2024subsurface, liu2023openillumination, zhang2025texverse, choi2025real]. In comparison, our work focuses on appearance capture of real-world objects. We categorize the existing object-centric datasets into several subgroups, as summarized in Tab. 1. The first group focuses on synthetic objects, which are highly scalable due to the procedural generation pipelines and the availability of public 3D models. ShapeNet-Intrinsics [shi2017learning], for example, constructs a large-scale object intrinsic dataset from 3D models in the ShapeNet [chang2015shapenet] collection. ABO [collins2022abo] introduces synthetic household objects with complex geometry and materials. Objaverse [deitke2023objaverse] and TexVerse [zhang2025texverse] significantly expand the diversity of objects by incorporating both synthetic and scanned real assets. While these datasets are large-scale, the quality of these datasets varies significantly across different objects, and a domain gap still remains between synthetic and real-world data. Consequently, effective curation strategies are required to filter out low-quality data, limiting their applicability in realistic 3D vision and relighting.
@@
Figure 2: (a) We visualize the statistics of OLATverse, including the material distribution and high-level object category distribution. (b) We also show comparison against OpenIllumination [liu2023openillumination] for the six largest material and object categories in terms of object count.
@@
The second group of datasets focuses on collecting real-world objects with high-fidelity appearance. Such datasets are typically collected either in specialized studios or in-the-wild environments. Among this group, DTC [dong2025digital] and OmniObject3D [wu2023omniobject3d] provide the meshes of thousands of real objects via 3D scanning. Although these two datasets are large-scale, they offer limited support for appearance analysis and lack control over lighting conditions. Specifically, OmniObject3D does not support appearance capture, and DTC only contains only 50 objects captured under two environmental illuminations. In addition, the material maps provided in DTC are manually created by artists to approximate the realistic appearance, which introduces a gap to the real-world appearance.
@@
The third group of datasets [bi2024gs3, liu2023openillumination, dihlmann2024subsurface] captures OLATs in studios with a well-designed lighting setups (e.g., lightstage [debevec2000acquiring], LCD, smartphone flash, etc.), facilitating detailed analysis of light transport. However, due to the complexity of the setup and capture procedure, these datasets remain small in scale and limited in diversity. The largest OLAT dataset, OpenIllumination [liu2023openillumination], contains only $64$ objects, limiting its application for comprehensive benchmarking and generative model training. In comparison with existing object-centric datasets, OLATverse encompasses $765$ real-world objects spanning diverse material categories and provides high-fidelity appearance under precisely controlled lighting, thereby achieving a combination of scale, diversity, and realism.
@@ 2.2 Inverse Rendering &\& Relighting
The goal of inverse rendering is to recover the intrinsic properties of scenes or objects (e.g., albedo, normal, roughness, and metallicity) to reproduce plausible renderings under novel illuminations. Traditional approaches typically leverage advanced differentiable renderers [nimier2019mitsuba, ravi2020accelerating, Laine2020diffrast] under reliable constraints provided by geometry to jointly optimize intrinsic properties and lighting by minimizing the discrepancy between input and rendered images. With the advances in deep learning, Neural Radiance Fields (NeRF)[mildenhall2021nerf] is introduced to jointly encode geometry and appearance in a single multi-layer perceptron (MLP), producing photorealistic rendering via volume rendering. NeRF demonstrates superior performance but suffers from slow rendering speed. To address this limitation, 3D Gaussian Splatting (3DGS)[kerbl20233d] is proposed to represent scenes with localized Gaussian kernels and can offer real-time efficiency through an efficient differentiable splatting process. Leveraging the benefits of these two representations, many recent inverse rendering methods [zhang2021nerfactor, shi2025gir, gao2024relightable, liang2024gs, du2024gs, zhang2024prtgaussian, liu2024bigs, jin2023tensoir, fan2025rng] extend NeRF or 3DGS with additional optimizable Bidirectional Reflectance Distribution Function (BRDF) modules and a differentiable renderer to recover intrinsic properties and lighting from multi-view input images. More recently, with the rapid progress of diffusion-based generative models, a new branch of research leverages the strong priors of pretrained diffusion models for inverse rendering [zeng2024rgb, ye2024stablenormal, liang2025diffusion, chen2024intrinsicanything] and relighting [jin2024neural, zeng2024dilightnet] tasks. These methods typically fine-tune pretrained diffusion models on datasets annotated with material maps or renderings under novel illuminations. During inference, the diffusion model extracts material reflectance properties or relit images through inverse denoising steps. In this work, we demonstrate that OLATverse serves as a comprehensive benchmark for object-centric inverse rendering and relighting tasks, and we evaluate several representative approaches as baselines on our selected validation dataset.
@@ 3.1 Overview
OLATverse comprises $765$ real-world objects with high diversity in material types, object categories, and physical sizes. As shown in (a) of Fig. 2, it covers over $13$ material categories, including wood, stone, plaster, fabric, plastic, metal, food, plant, ceramics, leather, wax, rubber, and paper. Furthermore, OLATverse consists of a wide range of common and uncommon real-world objects, spanning over $18.5\%$ of LVIS categories [gupta2019lvis], substantially outperforming existing real-world object datasets, such as OmniObject3D ($10.8\%$[wu2023omniobject3d], OpenIllumination ($4\sim 5\%$[liu2023openillumination] and DTC ($3\%$[dong2025digital]. Additionally, OLATverse is not limited to object physical dimension, encompassing a wide range of sizes from $5\mathrm{cm}$ to $100\mathrm{cm}$, which is broader than OpenIllumination dataset ($10\sim 20\mathrm{cm}$). These attributes highlight the diversity and scale of OLATverse, establishing it as the first comprehensive OLAT dataset of real-world objects.
@@
We compare OLATverse against one representative OLAT real dataset, OpenIllumination [liu2023openillumination]. We visualize the number of objects in the six largest material and high-level object categories, as illustrated in Fig. 2 (b). While OpenIllumination covers the similar material and object categories as OLATverse, the scale of each category is significantly smaller than our dataset. In addition, OpenIllumination does not provide auxiliary resources such as surface normals and diffuse albedo, limiting its applications in the multi-modal tasks.
@@ 3.3 Data Processing
Acquiring accurate camera parameters consistently across large-scale $765$ objects in a studio setup is a non-trivial task. Directly applying feature-based calibration algorithms to different objects leads to unstable and ambiguous calibration quality since the real-world objects vary widely in physical dimension, texture richness, and material properties, which limit the accuracy and consistency of the camera calibration process. To address this challenge, we leverage the fixed camera configuration of the lightstage, and perform a calibration session every $20\sim 30$ regular capture sessions, demonstrated in (b) of Fig. 3. During calibration session, we lock down several fixed objects with rich texture and Lambertian surfaces as calibration references, and employ the feature-based algorithms implemented in Metashape [agisoft_metashape] to recover both intrinsic and extrinsic camera parameters. For subsequent capture sessions, the previously estimated camera parameters are reused. All calibration captures are conducted under uniform white illumination to ensure robust and consistent feature detection. We quantitatively evaluate the calibration accuracy by computing the mean reprojection error of triangulated keypoints across the reference objects, yielding an average error of $0.86$ pixels. The positions of individual lights are also provided within the same canonical system by measuring their physical location. We visualize the calibrated cameras, light sources, and the reconstructed mesh of one calibration reference object in  Fig. 5. As is shown, the reconstructed mesh qualitatively aligns with real captured frames, demonstrating the accuracy of calibration process.
@@
Unlike human-centric datasets [teufel2025humanolat], where Sapiens [kim2025sapiensid] is tailored specifically for human body segmentation, efficiently producing high-quality masks for large-scale objects in our capture setup remains a challenging task. A straightforward solution is to employ Segment-Anything (SAM) [kirillov2023segment] with multiple bounding-box prompts for instance segmentation, similar to OpenIllumination [liu2023openillumination]. However, this strategy is inefficient and difficult to scale for a large-scale dataset. To efficiently tackle this challenge, we develop a simple yet effective semi-automatic segmentation pipeline that combines background matting (bgMatting) [lin2021real], SAM [kirillov2023segment] and RMBG-2.0 [bria2023rmbg2]. Specifically, we capture a foreground image $\mathbf{I}_{\text{fg}}$, containing the object and stand, and a background image $\mathbf{I}_{\text{bg}}$ only for the stand, as is shown in (d) of Fig. 3. We empirically observe that bgMatting is capable of roughly separating the object from the stand, but fails to preserve fine details around the contour of objects. SAM, when guided by bounding boxes generated from bgMatting, can achieve a clean separation between the stand and the object but at the cost of mask quality. In comparison to the other two methods, RMBG-2.0 captures the most detailed contours but consistently treats the stand as part of the foreground, yielding a wrong object mask. Therefore, we leverage the strengths of all three segmentation strategies, and compute the final object mask $\mathbf{M}_{\text{obj}}$ as follows:
@@
Figure 7: We visualize the inverse rendering and novel view synthesis results of several baseline methods (Mitsuba [nimier2019mitsuba], GS3 [bi2024gs3],BiGS [liu2024bigs] and RNG [fan2025rng]) evaluated on our validation dataset. In this figure, we show relit objects from inference views and light directions.
@@
where $\mathbf{M}_{1}$, $\mathbf{M}_{2}$ and $\mathbf{M}_{3}$ represent bgMatting [lin2021real], SAM [kirillov2023segment] and RMBG-2.0 [bria2023rmbg2], respectively. $\mathbf{M}_{\text{stup}}$ represents the mask of the stand. In Eq. 1, case a is applied to lower camera views, where part of the objects may be occluded by the stand, while case b is used for all other views. Then we refine the intermediate results $\mathbf{M}_{\text{obj}}^{*}$ using morphological transformations and remove the disconnected components to obtain the final clean mask $\mathbf{M}_{\text{obj}}$. This pipeline achieves a success rate of $95\%$ across all objects from all views in our setup. For the remaining failure cases, we create a lightweight user interface for manual correction. In  Fig. 6, we illustrate masks generated by three individual segmentation strategies and our method.
@@
In addition to diverse illuminations, OLATverse provides pseudo ground truth surface normals and diffuse albedo as auxiliary data to support multimodal tasks. We recover surface normals through photometric stereo solutions [guo2019relightables] by analyzing radiance variations of images captured under gradient illuminations. For all cameras, we captured images $\mathbf{I}_{\text{cg}}$ under color gradient illumination. To further cancel out the view-dependent specular reflection on the non-Lambertian surfaces [ma2007rapid, ghosh2009estimating], we apply linear polarized filters to five fixed cameras, and capture images under polarized gradient full bright illuminations, denoted as $\mathbf{I}_{\perp}$. The diffuse albedo $\mathbf{D}$ and surface normals $\mathbf{N}$ are computed as follows: